OpenStudy (girlstudy):

A tennis ball is struck and departs from the racket of Prof. Jacome horizontally with a speed of 30m/s. The ball hits the court at a horizontal distance of 20 m from the racket. How far above the court is the tennis ball when it leaves the racket?

OpenStudy (solomonzelman):

The tennis ball is travelling horizontally with a speed of 30m/s, and ends up 20m away from the racket. This is all we are given.

OpenStudy (solomonzelman):

So, we will assume the racket is \(y_0\) from the ground.

OpenStudy (solomonzelman):

\(y_0{~}-{\rm the~initial~hieght~of~the~ball}\). \(v_{x0}=v_x=30~{\rm m/s}\) (\(v_{x0}=v_x\) because air resistance is neglected) \( x=20\) (distance before landing)

OpenStudy (solomonzelman):

Can you work out the time it takes the ball to travel 20m horizontally?

OpenStudy (solomonzelman):

(You velocity in the horizontal is unchanged, because we neglect air resistance)

OpenStudy (girlstudy):

For time, can we use t= v - vo / a

OpenStudy (solomonzelman):

No, you don't need any kinematics for this.

OpenStudy (solomonzelman):

Your velocity in the horizontal is 30m/s (and it doesn't change). The distance in the horizontal that you traveled is 20m. So the time is?

OpenStudy (solomonzelman):

(Yes, it is that simple, because the motion in the horizontal and the motion in the vertical are independent of each other.)

OpenStudy (solomonzelman):

\(\displaystyle \frac{\Delta x}{\Delta t}=v_x\) Right? So, \(\displaystyle \Delta t=\frac{\Delta x}{v_x}=\frac{20{\rm m}}{30{\rm m/s}}=?\)

OpenStudy (girlstudy):

.666 seconds

OpenStudy (solomonzelman):

yes, 2/3 seconds.

OpenStudy (solomonzelman):

(I will use a fraction. I like to use exact values, and as I find the answer, then approximate .... well, in any case, we want the concept, so if you have any conceptual questions, let me know.)

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle v_{y}(t)=v_{y_0}-gt }\) familiar with this?

OpenStudy (girlstudy):

hmm is what is vy and vy0 in your formula?

OpenStudy (solomonzelman):

Alternatively, \(\color{black}{\displaystyle v=v_{0}-gt }\), and this is for the velocity in the y direction, where g is gravity.

OpenStudy (solomonzelman):

I am using a subscript of y to denote that it is a motion or velocity in the vertical.

OpenStudy (solomonzelman):

In other words, you agree that we can model the velocity in the\(y\) at any time \(t\) with \(\color{black}{\displaystyle v(t)=v_{0}-gt }\)

OpenStudy (solomonzelman):

and the consequence of this equation (via integration) is \(\color{black}{\displaystyle y(t)=(v_{0})t-\frac{1}{2}gt^2+y_0 }\)

OpenStudy (solomonzelman):

This equation models the vertical displacement. Seen this one before?

OpenStudy (girlstudy):

hmm well if your talking about integration i am not familiar with it. we're still just learning basic physics

OpenStudy (solomonzelman):

No, the equation, \(\color{black}{\displaystyle y(t)=(v_{0})t-\frac{1}{2}gt^2+y_0 }\)

OpenStudy (solomonzelman):

Did you see this (above) equation?

OpenStudy (solomonzelman):

Is this the first time you see this equation, or is it familiar to you?

OpenStudy (girlstudy):

i dont think ive seen this equation

OpenStudy (solomonzelman):

Oh, it's one of the kinematics, and this is the appropriate equation to use in this case.

OpenStudy (solomonzelman):

Why? Because you are given the time it takes to fall, t=2/3 seconds, you are given y(t)=0 at this time (b/c at this time ball hits the ground). Also you know gravity is 9.8m/s^2, and you know that the initial velocity in the vertical (the \(v_0\)) is 0.

OpenStudy (girlstudy):

Is this formula the same as y= V0yt+ 1/2 ayt ?

OpenStudy (solomonzelman):

yes, it is the same formula. \(\color{black}{\displaystyle y=v_{0y}+\frac{1}{2}a_yt}\)

OpenStudy (solomonzelman):

Do you know all of the information to plug in, or should we work on this together?

OpenStudy (solomonzelman):

(You are certainly welcome to try it, and I will check you .... wanna do this?)

OpenStudy (girlstudy):

Yes! i would like to do so

OpenStudy (solomonzelman):

Sure:) Go ahead, and take your time.

OpenStudy (girlstudy):

How would i go about finding Voy?

OpenStudy (solomonzelman):

First, I want to adjust your formula.

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle y(t)=v_{0y}+\frac{1}{2}a_yt+y_0}\)

OpenStudy (solomonzelman):

This is more appropriate, because the initial displacement in the y, is not 0. And y(t), because this is the height at time t.

OpenStudy (solomonzelman):

As far as the velocity goes, you are given \(v_{0x}\), and \(v_{0y}\) is not mentioned, so we are assuming that initial the ball's speed is perfectly horizontal.

OpenStudy (solomonzelman):

(In other words, your initial velocity in the y, just the instant the ball leave the racket, is equal to 0.)

OpenStudy (solomonzelman):

and knowing this you can just cross \(\color{black}{\displaystyle y(t)=\color{red}{\bcancel{\cancel{\color{black}{v_{0y}}}}}+\frac{1}{2}a_yt+y_0}\)

OpenStudy (solomonzelman):

So, with a zero initial velocity in the y you will have \(\color{black}{\displaystyle y(t)=\frac{1}{2}a_yt+y_0}\)

OpenStudy (solomonzelman):

The time it takes to fall is \(t=2/3{~}{\rm s}\), and at this time you know that \(y(t)=0\) (because this is when the ball reaches the ground). We also know \(\color{black}{\displaystyle a_y=g=-9.8~{\rm m/s^2}}\).

OpenStudy (solomonzelman):

So, there you have it all to find the initial displacement in the vertical, \(y_0\). (This \(y_0\) is the initial height of the ball, and this is the same height at the very instant the ball leaves the racket.)

OpenStudy (girlstudy):

oh so the Yo would be 20m ?

OpenStudy (solomonzelman):

Please read the question, and tell me what this value (20 m) corresponds to.

OpenStudy (solomonzelman):

You are given everything (I listed) ... all that remains for you is to plug in the information into \(\color{black}{\displaystyle y(t)=\frac{1}{2}a_yt+y_0 }\) and solve for \(y_0\).

OpenStudy (girlstudy):

Im not sure if i follow you

OpenStudy (girlstudy):

Its very confusing

OpenStudy (solomonzelman):

Do you understand why it is \(\color{black}{\displaystyle y(t)=\frac{1}{2}a_yt+y_0 }\), (without adding \(v_{0y}t\) to this)?

OpenStudy (solomonzelman):

(Because originally, the equation is \(\color{black}{\displaystyle y(t)=v_{0y}t+\frac{1}{2}a_yt+y_0 }\), however we got rid of that term by noticing something. -- What did we notice?)

OpenStudy (girlstudy):

Voy is not mentioned in the problem

OpenStudy (solomonzelman):

We noticed that \(\color{black}{\displaystyle v_{0t}=0 }\), because the only velocity that is there, is the velocity in the horizontal (and not in vertical).

OpenStudy (solomonzelman):

And since \(v_{0y}=0\) we can just write the same without \(v_{0y}t\).

OpenStudy (solomonzelman):

For this reason, instead of the original equation, I have now this: \(\color{black}{\displaystyle y(t)=\frac{1}{2}a_yt+y_0 }\)

OpenStudy (solomonzelman):

Can you tell me, what is \(\color{black}{\displaystyle a_y }\) ?

OpenStudy (girlstudy):

Oooh! ok ok Ay would be the acceleration

OpenStudy (solomonzelman):

Yes, and in this case the acceleration in the vertical is the what?

OpenStudy (solomonzelman):

BTW, COrrection t is squared. \(\color{black}{\displaystyle y(t)=\frac{1}{2}a_yt^2+y_0 }\)

OpenStudy (jonm111):